The calcium line of wavelength λ = 422.673 nm (P → S) exhibits normal Zeeman
splitting when placed in uniform magnetic field of 4 webers/meter2
. Calculate the
wavelength of three components of normal Zeeman pattern and the separation between
them
1
Expert's answer
2021-08-23T12:34:24-0400

Splitting between energy level in the presence of magnetic field is given by

"\frac{hc}{λ^2}Δλ=±\mu_BB \\

\mu_B = 9.27 \times 10^{-27} \;A \cdot m^2 \\

B=4 \;Wb/m^2 \\

Δλ = ± \frac{\mu_B B λ^2}{hc} \\

λ = 422.673 \; nm \\

Δλ = ± \frac{9.27 \times 4 \times (422.673 \times 10^{-9})^2}{6.626 \times 3 \times 10^{-26}} \\

= ± 3.332 \times 10^{-11} \\

= ± 0.033 \;nm"

The wavelength of three components of normal Zeeman pattern

λ, λ+Δλ, λ-Δλ = 422.673 nm, 722.706 nm, 422.640 nm

Splitting

Δλ=0.033 nm